8.        Random sampling from two normal populations produced the following results:

x1 = 412       s1 = 128      n1 = 150

x2 = 405     s2 = 54       n2 = 150

(a.) Can we infer at the 5% significant level that m1

is greater than m2 ?

(b.) Repeat part (a) decreasing the standard deviations to: s1 = 31 and s2 = 16

 

 

9.            In random samples of 25 from eah of two normal populations we found the following statistics:

x1 = 524      s1 = 129

x2 = 469     s2 = 141

(a.) Estimate the difference between the two population means with 95% confidence (b.) Repeat part (a) increasing the standard deviations

to: s1 = 255 and s2 = 260


1
Expert's answer
2021-11-10T18:11:55-0500

8 (a) Null hypothesis: "m{\scriptscriptstyle 1}≤m{\scriptscriptstyle 2}"

Alternative hypothesis: "m{\scriptscriptstyle 1}>m{\scriptscriptstyle 2}"

Since both sample sizes are greater than 30, we can use Z-score

According to the form of the null hypoyhesis, we should use right-tailed test

"P(Z>Z{\scriptscriptstyle cr})=0.05\implies Z{\scriptscriptstyle cr}=1.645"


The test value is:

"Z{\scriptscriptstyle test}={\frac {(x{\scriptscriptstyle 1}-x{\scriptscriptstyle 2})} {\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}}}={\frac {412-405} {\sqrt{{\frac {128^2} {150}}+{\frac {54^2} {150}}}}}={\frac 7 {11.3}}=0.62"

There is not enoughg evidencve to reject null hypothesis on th 0.05 significance level, so m1

is not greater than m2


(b) Now s1 = 31 and s2 = 16, which means our test statistic would be

"Z{\scriptscriptstyle test}={\frac {(x{\scriptscriptstyle 1}-x{\scriptscriptstyle 2})} {\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}}}={\frac {412-405} {\sqrt{{\frac {31^2} {150}}+{\frac {16^2} {150}}}}}={\frac 7 {2.85}}=2.456"

There is enoughg evidencve to reject null hypothesis on th 0.05 significance level, so m1

is greater than m2


9 (a) The difference between two sample means with 95% confidence is

"a \in (""(x{\scriptscriptstyle 1} - x{\scriptscriptstyle 2})-Cr{\scriptscriptstyle 0.95}*{\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}};(x{\scriptscriptstyle 1} - x{\scriptscriptstyle 2})+Cr{\scriptscriptstyle 0.95}*{\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}})"

Since we have small sample sizes, then it is appropriate to use t-statistic(two-tailed) with n - 1 degrees of freedom

"P(T(24)>Cr{\scriptscriptstyle 0.95})=0.025\implies Cr = 2.063"

"(x{\scriptscriptstyle 1} - x{\scriptscriptstyle 2})*Cr{\scriptscriptstyle 0.95}*{\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}})=(55-2.063*38.2;55+2.063*38.2)=(-23.8;133.8)"

(-23.8;133.8) is the 95% confidence interval


(b) for s1 = 255 and s2 = 260 we have

"(x{\scriptscriptstyle 1} - x{\scriptscriptstyle 2})*Cr{\scriptscriptstyle 0.95}*{\sqrt{({\frac {s^2{\scriptscriptstyle 1}} {n{\scriptscriptstyle 1}}} + {\frac {s^2{\scriptscriptstyle 2}} {n{\scriptscriptstyle 2}})}}})=(55-2.063*72.84;55+2.063*72.84)=(-95.27;205.29)"

(-95.27;205.29) is tha 95% confidence interval